Please help in solving the question  √7^log5 - √5^log7

Grade:9

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
given ques:√7^(log 5) -√5^(log 7)
=(7^(1/2))^(log 5) -( 5^(1/2))^(log 7)
=(7) ^( (1/2)log(5) ) -(5)^( (1/2)log(7))
=(7) ^ (log √5) - (5)^(log √7 )   
=(10^(log7))^(log5)-(10^(log 5))^(log 7)
=10 ^(log7 * log 5)-10^(log 5 * log 7)
=10^(log5 * log 7)-10^(log 5 * log7)
=0     (Ans.)
Hope it helped.
Thanks,
Soumya Ranjan Mohanty.